proofs For this assignment, know that: An integer is any countable number. Examples are: -3, 0,...

proofs

For this assignment, know that:
An integer is any countable number. Examples are: -3, 0, 5, 1337, etc.
A rational number is any number that can be written in the form a/b, a and b are integers in lowest terms, and b
cannot equal 0. Examples are 27, 22/7, -3921/2, etc.
A real number is any number that is not imaginary or infinity. Examples are 0, 4/3, square root of 2, pi, etc.

1. Prove or disprove: There exists no integers a and b such that 8a - 12b = 1
Answer #1

Equation is

8a-12b=1

4(2a-3b)=1

2a-3b=1/4

as the result of 2a-3b is >0 and a rational number, at least one of the a,b should be a rational number.

Ex:

a. a=2,b=1

2*2-3*1=1

b. a=-1, b=-3/4

-1*2 - -3*(-3/4) = 1/4

Hence proved. There exists no integers a and b such that 8a-12b=1